# Benchmarking Multimodal Retrieval Augmented Generation with Dynamic VQA Dataset and Self-adaptive Planning Agent

### v2

@@ Abstract
[…]
@@ 1 Introduction
Existing heuristic mRAG methods typically predefined fixed retrieval processes that ground all modalities into one primary modality (usually text), then retrieve for a single time. Despite the promising results, these retrieval strategies suffer from two issues: (1) Non-adaptive Retrieval Queries. These inflexible retrieval strategies fail to adapt to evolving contexts or intermediate findings within a question, hindering the model from re-retrieving to further comprehend, verify, or rethink the question. (2) Overloaded Retrieval Queries. These one-time retrieval strategies place too excessive burden on a single query, which contains multiple aspects. This potentially results in an influx of superficially relevant knowledge yet not essential to the question solving. Therefore, as shown in Figure 1, when faced with real-world questions requiring complex knowledge, current heuristic mRAG methods fail to provide sufficient and precise knowledge, due to their rigidity issues.
@@
To facilitate the mRAG research and bridge the dataset gap, we first construct a challenging dataset, Dyn-VQA, comprising 1,452 dynamic questions that require complex multimodal knowledge retrieval for solution. Dynamic questions are defined as questions that require the model to flexibly provide knowledge retrieval solutions, where the query, tool, and time of retrievals are all variable. These questions cannot be solved by a predefined retrieval process. Concretely, there types of dynamic questions are included in Dyn-VQA: (1) Questions with rapidly changing answers. Since the context knowledge of such question updates frequently, the retrieved content may be mixed with outdated and newer knowledge that is difficult to distinguish. This requires mARG methods to flexibly plan additional retrievals based on feedback from current retrieved content for further comprehension, rather than merely a one-time retrieval. (2) Questions requiring multi-modal knowledge. The knowledge necessary by Dyn-VQA spans various modalities. This demands that mRAG methods retrieve knowledge across diverse modalities with tailored retrieval APIs, differing from most VQA datasets limited in seeking textual knowledge with multimodal questions. (3) Multi-hop questions. Questions in Dyn-VQA necessitate varied reasoning hops for solution, which entails that mRAG methods conduct various retrieval steps. While existing VQA datasets primarily focus on two-hop question, i.e., identifying visual concepts via text and then answering single-hop textual question.
@@
To address these issues, we further propose a self-adaptive planning agent for multimodal retrieval, OmniSearch [^2] We hope OmniSearch can achieve Omnipotent Multimodal Search for solving real-world questions in the future.. The underlying idea is to emulate the human behavior in question solution which dynamically decomposes complex multimodal questions into sub-question chains with retrieval action. At each step, OmniSearch flexibly adjusts the next action according to question-solving state and retrieved content, with diverse purposes such as deepening comprehension of retrieved content, modifying retrieval method for current sub-question, proposing the next sub-question, etc. It is noteworthy that OmniSearch can serve as a plug-and-play RAG module, cooperating with arbitrary MLLMs to augment their capability in addressing complex dynamic questions. Two different versions of OmniSearch are developed based on closed-source GPT-4V (Achiam et al., 2023) and open-source Qwen-VL-Chat (Bai et al., 2023a), respectively. As far as we know, OmniSearch is the first multimodal retrieval agent for VQA tasks.
@@
  • We reveal that existing VQA-based mRAG benchmarks fail to reflect the feature that real-world questions require dynamic knowledge retrieval, and propose novel Dyn-VQA dataset, which contains three types of dynamic questions.
@@ 2.1 Multimodal Retrieval Augmented Generation
One category employs visual encoding model for direct image representation, and then retrieves the knowledge from knowledge base with the highest feature similarity. For instance, KAT (Gui et al., 2022) and Revive (Lin et al., 2022) both use the image encoder of CLIP (Radford et al., 2021) for retrieval. The other category (Hu et al., 2023; Yang et al., 2022; Lin et al., 2024) first transforms the input image into textual representation utilizing off-the-shelf tools and then performs text retrieval. For example, RA-VQA (Lin & Byrne, 2022) and RA-VQA-v2 (Lin et al., 2024) first use existing object detection and image captioning models to convert images to text, and than perform dense passage retrieval to fetch relevant text documents.
@@ 4.2 OmniSearch Framework
Heuristic mRAG methods encounter the issues posed by rigid retrieval processes. A more flexible model that can plan the retrieval strategy tailored to the specific question and dynamically adjusts it throughout the process is desired. Therefore, we propose a self-routing framework called OmniSearch for retrieval planning. The underlying idea behind is to mimic the human of incrementally decomposing the complex question into a sequence of solution actions. As illustrated in Figure 3, overall framework consists of three module. The planning agent is the core module that formulates sub-questions and plans the subsequent retrieval action based on real-world feedback (i.e., retrieved content or solver output). Actual retrieval action is execute by the retriever. Then, sub-question solver generates feedback on sub-question based on the retrieval content and update it to the planner.
@@ B.3 VQA datasets as mRAG Benchmarks
Knowledge-seeking VQA datasets (Marino et al., 2019; Jain et al., 2021; Schwenk et al., 2022) are widely employed to evaluate the performance of mRAGs, which rely on external information to address open-domain visual questions. For instance, the recently introduced Wikipedia-based VQA dataset, InfoSeek (Chen et al., 2023), emphasizes fine-grained entity knowledge for open-domain questions. A-OKVQA (Schwenk et al., 2022) is a new knowledge-based VQA benchmark that necessitates a broad spectrum of commonsense and world knowledge. As illustrated in Table 1, several knowledge-seeking VQA datasets have been proposed in recent years.
@@ C.3 Training Details
C.4 Prompt Templates Used in Methods
The prompts used for original MLLMs, heuristic mRAG-empowered MLLMs, and our OmniSearch are listed in Table 10. Prompts are differentiated between Chinese and English questions.
C.5 Supplement to Main Experiments
@@ D.1 Analysis Experiments on Retrieved Content
Table 10: The prompt used in original MLLMs, heuristic mRAG-empowered MLLMs, and our OmniSearch.
{CJK}
[…]

### v4

@@ Abstract
Multimodal Retrieval Augmented Generation (mRAG) plays an important role in mitigating the “hallucination” issue inherent in multimodal large language models (MLLMs). Although promising, existing heuristic mRAGs typically predefined fixed retrieval processes, which causes two issues: (1) Non-adaptive Retrieval Queries. (2) Overloaded Retrieval Queries. However, these flaws cannot be adequately reflected by current knowledge-seeking visual question answering (VQA) datasets, since the most required knowledge can be readily obtained with a standard two-step retrieval. To bridge the dataset gap, we first construct Dyn-VQA dataset, consisting of three types of “dynamic” questions, which require complex knowledge retrieval strategies variable in query, tool, and time: (1) Questions with rapidly changing answers. (2) Questions requiring multi-modal knowledge. (3) Multi-hop questions. Experiments on Dyn-VQA reveal that existing heuristic mRAGs struggle to provide sufficient and precisely relevant knowledge for dynamic questions due to their rigid retrieval processes. Hence, we further propose the first self-adaptive planning agent for multimodal retrieval, OmniSearch. The underlying idea is to emulate the human behavior in question solution which dynamically decomposes complex multimodal questions into sub-question chains with retrieval action. Extensive experiments [^1] This work was led by Xinyu Wang at Alibaba Group, with Yangning, Yinghui, and Zheng Zhang as his interns. The code and dataset is open-sourced at https://github.com/Alibaba-NLP/OmniSearch. prove the effectiveness of our OmniSearch, also provide direction for advancing mRAG.
@@ 1 Introduction
Existing heuristic mRAG methods typically predefined fixed retrieval processes that ground all modalities into one primary modality (usually text), then retrieve for a single time. Despite the promising results, these retrieval strategies suffer from two issues: (1) Non-adaptive Retrieval Queries refer to the fixed retrieval processes and query structures of heuristic mRAG methods. These inflexible retrieval strategies fail to adapt to evolving contexts or intermediate findings within a question, hindering the model from re-retrieving to further comprehend, verify, or rethink the question. For example, in Figure 2, question (a) asks, “What is his (Cillian Murphy’s) latest film?” A fixed retrieval process returns multiple relevant films, but heuristic methods fail to construct further retrieval based on the retrieved content to distinguish between the sequence of different films. (2) Overloaded Retrieval Queries refer to heuristic retrieval methods refer to heuristic mRAG methods merely format a single query by concatenating textual descriptions of objects in images with input questions. A single query carries multiple retrieval aspects, leading to ambiguous retrieval and influx of superficially relevant knowledge yet not essential to the question solving. For example, in Figure 2, question (c) asks, “Which one of them (two actors, Ling Jia, and Teng Shen) grossed more?” Heuristic methods might generate a single query like “Ling Jia, Teng Shen, Which one of them grossed more?”, which contains the intent to retrieve box office information for both actors. This mixed query conversely fails to provide precise knowledge for each individual aspect. Therefore, as shown in Figure 1, when faced with real-world questions requiring complex knowledge, current heuristic mRAG methods fail to provide sufficient and precise knowledge, due to their rigidity issues.
@@
To bridge the mRAG dataset gap, we first construct a challenging dataset, Dyn-VQA, comprising 1,452 dynamic questions that require complex multimodal knowledge retrieval for solution. Dynamic questions are defined as questions that require the model to flexibly provide knowledge retrieval solutions, where the query, tool, and time of retrievals are all variable. These questions cannot be solved by a predefined retrieval process. Concretely, there types of dynamic questions are included in Dyn-VQA: (1) Questions with rapidly changing answers. Since the context knowledge of such question updates frequently, the retrieved content may be mixed with outdated and newer knowledge that is difficult to distinguish. This requires mARG methods to flexibly plan additional retrievals based on feedback from current retrieved content for further comprehension, rather than merely a one-time retrieval. (2) Questions requiring multi-modal knowledge. The knowledge necessary by Dyn-VQA spans various modalities. This demands that mRAG methods retrieve knowledge across diverse modalities with tailored retrieval APIs, differing from most VQA datasets limited in seeking textual knowledge with multimodal questions. (3) Multi-hop questions. Questions in Dyn-VQA necessitate varied reasoning hops for solution, which entails that mRAG methods conduct various retrieval steps. While existing VQA datasets primarily focus on two-hop question, i.e., identifying visual concepts via text and then answering single-hop textual question.
@@
To address these issues, we further propose a self-adaptive planning agent for multimodal retrieval, OmniSearch [^2] We aim for OmniSearch to achieve Omnipotent Multimodal Search, solving real-world issues in future.. The underlying idea is to emulate the human behavior in question solution which dynamically decomposes complex multimodal questions into sub-question chains with retrieval action. At each step, OmniSearch flexibly adjusts the next action according to question-solving state and retrieved content, with diverse purposes such as deepening comprehension of retrieved content, modifying retrieval method for current sub-question, proposing the next sub-question, etc. It is noteworthy that OmniSearch can serve as a plug-and-play RAG module, cooperating with arbitrary MLLMs to augment their capability in addressing complex dynamic questions. Two different versions of OmniSearch are developed based on closed-source GPT-4V (Achiam et al., 2023) and open-source Qwen-VL-Chat (Bai et al., 2023a), respectively. As far as we know, OmniSearch is the first multimodal retrieval agent for VQA tasks with self-adaptive planning and scalable submodules.
@@
  • We reveal that current VQA-based mRAG benchmarks don’t reflect real-world needs for dynamic knowledge retrieval and propose the Dyn-VQA dataset with three types of dynamic questions.
@@ 2.1 Multimodal Retrieval Augmented Generation
One category employs visual encoding model for direct image representation, and then retrieves the knowledge from knowledge base with the highest feature similarity. For instance, KAT (Gui et al., 2022) and Revive (Lin et al., 2022) both use the image encoder of CLIP (Radford et al., 2021) for retrieval. The other category (Hu et al., 2023; Yang et al., 2022; Lin et al., 2024) first transforms the input image into textual representation utilizing off-the-shelf tools and then performs text retrieval. For example, RA-VQA (Lin & Byrne, 2022) and RA-VQA-v2 (Lin et al., 2024) first use existing object detection and image captioning models to convert images to text, and than perform dense passage retrieval to fetch relevant text documents. Several studies (Yao et al., 2023; Xu et al., 2023) have preliminarily explored into agentic RAG, but their primary focus was on the text domain. Chen et al. (2024a) proposed a causality-enhanced agent framework focus on unimodal biases in MLLMs, while it lacks plug-and-play capabilities.
@@ 4.2 OmniSearch Framework
Heuristic mRAG encounter the issues posed by rigid retrieval processes. A more flexible model that can plan the retrieval strategy tailored to the specific question and dynamically adjusts it throughout the process is desired. Therefore, we propose a self-routing framework called OmniSearch for retrieval planning. The underlying idea behind is to mimic the human of incrementally decomposing the complex question into a sequence of solution actions. As illustrated in Figure 3, overall framework consists of three module. The planning agent is the core module that formulates sub-questions and plans the subsequent retrieval action based on real-world feedback (i.e., retrieved content or solver output). Actual retrieval action is execute by the retriever. Then, sub-question solver generates feedback on sub-question based on the retrieval content and update it to the planner.
@@
The proposal of OmniSearch is inspired by Chain-of-Thought (CoT) Wei et al. (2022), but differs from it essentially. The fundamental distinction between OmniSearch, as a multimodal agent, and CoT is its ability to utilize tools, interact with the environment, and response to the environment Zhang et al. (2023). In contrast, CoT methods primarily stimulate the model’s inherent logical reasoning capabilities through prompts. The CoT approach is unable to decouple intermediate processes, therefore can not be integrated with retrieval tools.
@@ 7 Conclusion
Acknowledgements
This research project is supported by National Natural Science Foundation of China (Grant No.62276154), Research Center for ComputerNetwork (Shenzhen) Ministry of Education, the Natural Science Foundation of Guangdong Province (Grant No.2023A1515012914 and 440300241033100801770), Basic Research Fund of Shenzhen City (Grant No.JCYJ20210324120012033, JCYJ20240813112009013 and GJHZ20240218113603006), the Major Key Projectof PCL for Experiments and Applications (PCL2022A05 and PCL2023A09), NSF under Grants III-2106758, and POSE-2346158.
@@ B.3 VQA datasets as mRAG Benchmarks
Knowledge-seeking VQA datasets (Marino et al., 2019; Jain et al., 2021; Schwenk et al., 2022; Kil et al., 2024; Chen et al., 2024b) are widely employed to evaluate the performance of mRAGs, which rely on external information to address open-domain visual questions. For instance, the recently introduced Wikipedia-based VQA dataset, InfoSeek (Chen et al., 2023), emphasizes fine-grained entity knowledge for open-domain questions. A-OKVQA (Schwenk et al., 2022) is a new knowledge-based VQA benchmark that necessitates a broad spectrum of commonsense and world knowledge. As illustrated in Table 1, several knowledge-seeking VQA datasets have been proposed in recent years.
@@ C.3 Training Details
C.4 Supplement to Main Experiments
@@ D.2 Consistency of Different Evaluation Metrics
D.3 Supplementary Analysis of Computational Costs
In Table 13, we further reported the computational cost of OmniSearch on questions with different answer update frequencies (fast, slow, never). The results indicate that:
(1) Overall, OmniSearch consumes more tokens for more complex questions, such as those with fast or slow answer updates, primarily because these questions inherently require more retrieval steps.
(2) The difference in resource consumption for questions of varying difficulty is more pronounced for the smaller model OmniSearch (Q). This is due to the behavioral differences between GPT-4V and Qwen-VL. Specifically, as described in the Section 6.1, GPT-4V tends to be more rigorous in question-solving, proactively planning verification retrievals to validate final answers, leading to retrieval processes exceeding three steps even for some relatively easier questions.
(3) By comparing rows 1 and 2 in the Table 13, we can observe that after replacing OmniSearch (G)’s Sub-question Solver with a smaller Qwen-VL-Chat, the total token consumption of Qwen-VL-Chat and GPT-4V is stil comparable to the original model. This suggests that OmniSearch (G) effectively offloads computational burden to smaller models without a significant drop in overall performance.
Table 13: The token cost of OmniSearch on questions with different answer update frequencies.
Planning Model Sub-question Solver Fast Slow Never
Input T. Output T. Perf. Input T. Output T. Perf. Input T. Output T. Perf.
OmniSearch (G) OmniSearch (G) 3098.5 (G) 466.3 (G) 44.04 3036.5 (G) 477.0 (G) 49.58 2974.6 (G) 483.9 (G) 54.45
OmniSearch (G) Qwen-VL-Chat
1268.2 (G)
+ 2165.4 (Q)
403.1 (G)
+ 148.4 (Q)
38.65
1214.1 (G)
+ 1925.7 (Q)
358.5 (G)
+ 112.7 (Q)
44.68
1185.6 (G)
+ 2138.8 (Q)
398.2 (G)
+ 119.3 (Q)
52.25
OmniSearch (Q) OmniSearch (Q) 10258.2 (Q) 638.1 (Q) 35.16 9874.8 (Q) 634.2 (Q) 40.89 8866.9 (Q) 475.3 (Q) 45.52
OmniSearch (Q) GPT-4V
2508.8 (G)
+ 933.0 (Q)
269.3 (G)
+ 547.8 (Q)
37.14
2452.5 (G)
+ 1000.9 (Q)
233.6 (G)
+ 490.5 (Q)
42.82
2209.6 (G)
+ 1024 (Q)
329.7 (G)
+ 606.3 (Q)
47.48
Additionally, we also reported the average latency for each question in Table 14. The results show that substituting some modules in OmniSearch with smaller models effectively reduces latency. Moreover, the ratio of search time to model inference time is roughly 2:3, indicating significant potential for optimization in both aspects. It is important to note that latency is a complex system engineering issue that involves not only the model’s complexity but also factors network configuration of search APIs, caching strategies for retrieval content, inference model acceleration, and hardware FLOPS, etc.
Table 14: The average model latency for question in Dyn-VQA.
Planning Model Sub-question Solver Search Time (s/%) Inference Time (s/%) Total Time (s)
OmniSearch (G) OmniSearch (G) 14.1 (44.8%) 18.4 (55.2%) 31.5
OmniSearch (G) Qwen-VL-Chat 12.3 (44.6%) 15.3 (55.4%) 27.6
OmniSearch (Q) OmniSearch (Q) 8.5 (38.3%) 13.7 (61.7%) 22.2
OmniSearch (Q) GPT-4V 11.8 (45.0%) 14.4 (55.0%) 26.2